OpenStudy (anonymous):

Solve by elimination method: 9x-9y=36 7y-3x=-14 Another problem: 0.3x-0.2y=4 0.4x+0.5y=1

OpenStudy (anonymous):

Hello i will help you today

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

.3x - .2y = 4 -->(1) .4x + .5y = 1 -->(2) (1)x5 you get 1.5x - y = 20 (2)x2 you get .8x + y = 2 Add up these two, you get 2.3x = 22 x = 220/23 Substitute x back into any equation, you get y = -130/23 and 5x +3y = -11 -- (1) 7x - 2y =19 -- (2) (1)x2 you get 10x + 6y = -22 (2)x3 you get 21x - 6y = 57 Add the two, you get 31x = 35 So x = 35/31 Substitute x back into any equation, you get y = 172/31 and 9x - 9y =36 -- (1) 7y - 3x = -14 -- (2) (2)x3 you get -9x + 21y = -42 Add this to (1), you get 12y = -6, so y = -1/2 Substitute y back into any equation, you get x = 7/2

OpenStudy (anonymous):

tell me if that helps at all

OpenStudy (anonymous):

the one problem i think you wrote wrong it was supposed to be 5x+2y=-7 7x-3y=11

OpenStudy (anonymous):

ok i am srry hold on

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

wait nevermind i wrote it wrong

OpenStudy (anonymous):

5x+2y=-13 7x-3y=17

OpenStudy (anonymous):

ok here we go 9x-9y=36 => x - y = 4 => x = 4+y 7y-3x=-14 => 7y - 3(4+y) = -14 7y - 12 - 3y = -14 4y = -2 y = -1/2 x = 4-1/2 = 7/2 (x,y) = (7/2, -1/2)

OpenStudy (anonymous):

would you mind helping me on a few more

OpenStudy (anonymous):

no i would love to

OpenStudy (anonymous):

5x+2y=-13 7x-3y=17 0.05x+0.25=11 0.15x+0.05y=12

OpenStudy (anonymous):

Substitution method: 7m+n=46 m-3n=16

OpenStudy (anonymous):

Im really having trouble with the substitution one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

umm i cant to seem to get this won hold on and i am going to have some won help if that is ok with you

OpenStudy (anonymous):

well what would the elimination problems be i sent you

OpenStudy (anonymous):

ok is this it tight here Start by multiplying the first equation by 2, and the second by 3: 10x + 6y = -22 21x - 6y = 39 Now add the two together, you y's cancel out , "eliminating" it. 21x = 17 x = 17/21 5x + 3y = -11 5(17/21) + 3y = -11 85/21 + 3y = -11 85 + 63y = -231 63y = -146 y = -146 / 63

OpenStudy (anonymous):

okay thanks how about 0.05x+0.25y=11 0.15x+0.05y=12

OpenStudy (anonymous):

ok here we go 5x + 25 y = 1100 15 x + 5 y = 1200 - 15x - 75y = - 3300 15x + 5y = 1200-----------ADD - 70 y = - 2100 y = 30 5x + 750 = 1100 5x = 350 x = 70 x = 70 , y = 30

OpenStudy (anonymous):

wow thanks your like a lifesaver haha

OpenStudy (anonymous):

np i love to help

OpenStudy (anonymous):

5x+5y=-11 7x-3y=17

OpenStudy (anonymous):

this is done by elimination

OpenStudy (anonymous):

i am on it

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

5x+5y=-11 multiply by 2 ==> 10x + 10y = -22 7x-2y=17 multiply by 5 ===> 35x - 10y = 85 Add the two equations: 10x + 35x = -22+85 45x = 63 x = 63/45 = 7/5 substitute into original equation: 5(7/5) + 5y = -11 5y = -18 y = -18/5 Solution set: (x,y) = (7/5, -18/5) ok do this help

OpenStudy (anonymous):

it told me i was wrong

OpenStudy (anonymous):

ummm that is weird

OpenStudy (anonymous):

what r ypu working on so i know i might know why

OpenStudy (anonymous):

its a math program and I have to enter the answers in and it lets me know if im wrong or not then submits it to my teacheer part of my homeschool

OpenStudy (anonymous):

ok i do home schooling to i think i know what programe you talking about

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

umm that i weird i woundr were we went wrong

OpenStudy (anonymous):

i dunno im trying to figure it out too

OpenStudy (anonymous):

Ill rewrite the problem

OpenStudy (anonymous):

5x+5y=-11 7x-3y=17

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok i think i got it Let's just multiply everything by 100 to eliminate the decimals. 5x + 25y = 1100 15x + 5y = 1200 Multiply the first equation by 3 15x + 75y = 3300 15x + 5y = 1200 Subtract the second equation from the first 70y = 2100 7y = 210 y = 30 15x + 5(30) = 1200 15x + 150 = 1200 15x = 1200 -150 15x = 1050 x = 1050/15 = 70 Let's check in the original equations. 0.05(70) + 0.25(30) = 3.50 + 7.50 = 11 check. 0.15(70) + 0.05(30) = 10.50 + 1.50 = 12 check. 8 months ago Report Abuse

OpenStudy (anonymous):

srry about the 8 month ago i did that srry

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

i was working on something els and it got out up there

OpenStudy (anonymous):

did you see the word problem i posted could you help me with that

OpenStudy (anonymous):

did you ever figure out that one problem for me